# Gaussian process regression + deep neural network autoencoder for probabilistic surrogate modeling in nonlinear mechanics of solids

### v1

@@ 3.3 Predictions of the GP+autoencoder framework
(c) Absolute error
(d) Reconstruction error
(e) GP component error
(f) Uncertainty prediction
Figure 7: Deformation of the 2D beam when applied with the point load on the top right corner node. Nodal displacements obtained using the (a) proposed framework (b) FEM respectively. (c) Nodal error and (d) reconstruction error of the framework. (e) GP component error. (f) Uncertainty predicted by the framework for the full field solution.
Similarly, we analyse the test example of 3D liver case with the maximum nodal displacement of 5.74 m. Figure 8 shows that the framework is able to accurately predict the deformation responses of the 3D liver geometry subjected to external body forces. Figure 8a, 8b show that the nodal displacements predicted by the framework are in a great agreement with the FEM solution. Figure 8c, 8d present the framework and reconstruction errors which are extremely low compared to the magnitude of displacements. For the maximum displacement node, the prediction error is $0.2\%$ of its magnitude. Directing attention to Figure 8e and Figure 8f, these figures present the error contribution of the GP component and the predicted uncertainty respectively. Again, the nodal GP errors are within the respective nodal uncertainty values thereby indicating that the framework is able to make reliable predictions.
Overall in addition to the framework error, its breakdown into reconstruction and GP error provides a clear and accessible understanding of how each component contributes to the overall performance of the framework.
@@
(c) Absolute error
(d) Reconstruction error
(e) GP component error
(f) Uncertainty prediction
Figure 8: Deformation of the 3D liver for the highest nodal displacement case, subjected to external body force. Nodal displacements obtained using the (a) proposed framework (b) FEM respectively. (c) Nodal error and (d) reconstruction error of the framework. (e) GP component error. (f) Uncertainty predicted by the framework for the full field solution.
3.4 Overall accuracy over the test sets
The performance of the framework over the entire test set is computed using the mean absolute error metric. For the $m^{th}$ test example, the mean error is given as follows:
$e_{m}=\frac{1}{\mathcal{F}}\sum_{i=1}^{\mathcal{F}}{|\tilde{{\bf u}}^{*}_{\mu}({\bf f}_{m})^{i}-{\bf u}_{m}^{i}|}.$ (22)
$\mathcal{F}$ is the number of DOFs of the mesh representing the full field space, $\tilde{{\bf u}}^{*}_{\mu}({\bf f}_{m})$ is the mean prediction of the GP+Decoder framework (full field prediction) as described through Eq.(14) and ${\bf u}_{m}$ is the finite element solution for the full field space. To have a single validation metric over the entire test set, we compute the average mean norm $\bar{e}$ and the corrected sample standard deviation $\sigma(e)$ as follows:
@@ 3.4 Overall accuracy over the test sets
The proposed framework predicts probabilistic displacement fields corresponding to the full field mesh. First, we analyse the prediction performance by comparing mean predictions of the framework to the true FEM solutions. Table 2 presents error metrics for both benchmark cases. Additionally, the maximum nodal displacements for both cases are also provided, i.e., we compute displacements of all the nodes for every single example, and provide the maximum nodal displacement in the table. Table 2 shows that both mean and maximum errors are fairly low when compared to the displacement magnitudes. The proposed framework is capable of predicting mechanical deformation responses with very low errors.
@@
3D liver 400 0.9 E-3 0.9 E-3 7.7 E-2 5.74
Table 2: Error metrics for 2D and 3D test sets for predictions using the proposed GP+autoencoder framework. $M$ stands for the number of test examples, and $\bar{e}$, $\sigma(e)$ and $e_{\text{max}}$ are error metrics defined in Section LABEL:sec:_validation_metric.
Further, we analyze test examples with the maximum nodal displacement for both benchmark cases. In both cases, we plot nodal error contours representing the absolute difference between the mean displacement predicted by the framework and the true FEM solution. We also plot uncertainty estimates obtained using the proposed framework (see Eq.(14)).
3.5 Latent vs. reconstructed uncertainties
GPs are utilized to predict outcomes in the latent space. The GP forecasts a Gaussian probability distribution for each latent DOF. To ensure the efficiency and accuracy of GP’s predictions, we provide additional accuracy measures. First, we define prediction of a particular DOF as healthy, if the corresponding true solution lies within two standard deviations of the mean value.
@@
Figure 9: Latent distribution prediction for two randomly chosen test examples from the liver test dataset. All the 16 true latent values are transformed to an uniform scale, along with the displacement random variables. True latents lying within the two standard deviations (±$\pm 2\sigma$ region represented by shaded gray region) are indicated with green lines, while the latents lying outside are indicated with red colors. (a) All true latent displacements lie within ±$\pm\sigma$, which we classify as a correct prediction. (b) Three true latent values don’t lie within ±$\pm 2\sigma$, which we classify as an incorrect prediction.
@@ 3.6 Uncertainty for missing data region
In the second case, we use a new test dataset in which two dimensional input forces are randomly generated as denoted by red dots in Figure 10b. In Figure 11b we plot the uncertainty for the first latent DOF. It is evident that the uncertainty shows an increasing trend in the region not supported by training data. All the 20 latent DOFs show a very similar trend.
Elevated uncertainty levels are clearly elevated for areas of sparse training data. In the presented examples, these uncertainties correlate with the prediction errors, which can provide useful information. However, it’s crucial to emphasize that the predictive capabilities in the data-sparse and extrapolated regions can never be assured, with the tendency to deteriorate with the distance from the training data.
@@ 4 Conclusions
This work presents a novel approach for probabilistic surrogate modeling of high-fidelity simulations. It combines Gaussian processes with reduced-order modeling to efficiently simulate the mechanics of solids while accounting for uncertainties. The reduced representations of high-dimensional displacement data are achieved using autoencoder neural networks, allowing for drastic non-linear compression. For instance, this method reduced approximately 9100 degrees of freedom (DOFs) to 16 latent-space components, as demonstrated with the liver example. This strategy significantly reduces the computational cost of probabilistic predictions using Gaussian processes by lowering the output dimensionality, which is crucial because this cost can scale linearly to cubically with the output dimension size. We demonstrate that the proposed framework can accurately predict non-linear hyper-elastic deformations of solid bodies, along with the associated model uncertainties.
@@
This study opens up several other opportunities for future research. One exciting direction is applying this approach to probabilistic simulations of time and path dependent problems. One could use latent state to track how solutions evolve in such cases, which would make the whole process much faster and more efficient [63]. Another direction would be to exploit state of the art graph autoencoder networks to find compressed representations for arbitrary high-dimensional meshes [64]. This advancement would overcome the limitations of fully connected networks, which are recognized for their challenges in handling problems with a high number of dimensions. Finally, we believe that the proposed approach can find direct application in research and development, which is not restricted to problems in mechanics but covers a broader spectrum of engineering and scientific domains.

### v2

@@ 3.4 Performance study of the GP+autoencoder framework
(c) Absolute error, $e_{\text{f}}$, see Eq. (20)
(d) Reconstruction error, $e_{r}$, see Eq. (21)
(e) GP component error, $e_{\text{GP}}$, see Eq. (22)
(f) Uncertainty prediction, $(u_{\sigma}^{*})^{2}$, see Eq. (14)
Figure 7: Deformation of the 2D beam subjected to the point load (-2.00, 2.19) N on the top right corner node. Nodal displacements obtained using the (a) proposed framework (b) FEM respectively. (c) Nodal error and (d) reconstruction error of the framework. (e) GP component error. (f) Uncertainty predicted by the framework for the full field solution.
For the 2D beam example, we analyse the test example in with the maximum nodal displacement of 1.16 m, see Figure 7. The prediction of displacement field obtained using the proposed framework which is in a very good agreement with the FEM solution, as presented in the Figures 7a-b. Figure 7c provides nodal error of the entire framework. While Figure 7d and Figure 7e provide reconstruction error and GP error components as described by Eq.(21) and Eq.(22) respectively. Error plots show that all the errors are fairly low when compared to the magnitude of the displacement solution thereby displaying the capability of the framework in making accurate predictions. Figure 7f provides nodal uncertainty estimates predicted by the framework (2 standard deviations), as obtained from Eq.(14). One can observe that the GP component errors are within the uncertainty bounds for respective nodal predictions, i.e., in other words, true solution of the framework is contained within the displacement distribution for all the nodes.
@@
(c) Absolute error, $e_{\text{f}}$, see Eq. (20)
(d) Reconstruction error, $e_{r}$, see Eq. (21)
(e) GP component error, $e_{\text{GP}}$, see Eq. (22)
(f) Uncertainty prediction, $(u_{\sigma}^{*})^{2}$, see Eq. (14)
Figure 8: Deformation of the 3D liver for the highest nodal displacement case, subjected to external body force (0.71, 0.69, -0.78) N/kg. Nodal displacements obtained using the (a) proposed framework (b) FEM respectively. (c) Nodal error and (d) reconstruction error of the framework. (e) GP component error. (f) Uncertainty predicted by the framework for the full field solution.
Similarly, we analyse the test example of 3D liver case with the maximum nodal displacement of 0.14 m. Figure 8 shows that the framework is able to accurately predict the deformation responses of the 3D liver geometry subjected to external body forces. Figure 8a, 8b show that the nodal displacements predicted by the framework are in a great agreement with the FEM solution. Figure 8c, 8d present the framework and reconstruction errors which are extremely low compared to the magnitude of displacements. For the maximum displacement node, the prediction error is $0.2\%$ of its magnitude. Directing attention to Figure 8e and Figure 8f, these figures present the error contribution of the GP component and the predicted uncertainty respectively. Again, the nodal GP errors are within the respective nodal uncertainty values thereby indicating that the framework is able to make reliable predictions.
The performance of the framework over the entire test set is computed using the mean absolute error metric. For the $m^{\text{th}}$ test case, the mean error is given as follows:
$e_{m}=\frac{1}{\mathcal{F}}\sum_{i=1}^{\mathcal{F}}{|\tilde{{\bf u}}^{*}_{\mu}({\bf f}_{m})^{i}-{\bf u}_{m}^{i}|},$ (23)
where $\mathcal{F}$ is the number of DOFs of the mesh representing the full field space, $\tilde{{\bf u}}^{*}_{\mu}({\bf f}_{m})$ is the mean prediction of the GP+Decoder framework (full field prediction) as described through Eq.(14) and ${\bf u}_{m}$ is the finite element solution for the full field space. To have a single validation metric over the entire test set, we compute the average mean norm $\bar{e}$ and the corrected sample standard deviation $\sigma(e)$ as follows:
@@
The proposed framework predicts probabilistic displacement fields corresponding to the full field mesh. First, we analyse the prediction performance by comparing mean predictions of the framework to the true FEM solutions. Table 2 presents error metrics for both benchmark examples. Additionally, the maximum nodal displacements for both examples are also provided, i.e., we compute displacements of all the nodes for every single test case, and provide the maximum nodal displacement in the table. Table 2 shows that both mean and maximum errors are fairly low when compared to the displacement magnitudes, which demonstrates that the proposed framework is capable of predicting mechanical deformation responses with very low errors.
@@
3D liver 400 0.9 E-4 0.3 E-4 2.3 E-3 0.14
Table 2: Error metrics for 2D and 3D test sets for predictions using the proposed GP+autoencoder framework. $M$ stands for the number of test examples, and $\bar{e}$, $\sigma(e)$ and $e_{\text{max}}$ are error metrics defined in Equations (24)-(25).
3.5 GP performance in the latent space
GPs are utilized to predict outcomes in the latent space. The indpendent GPs forecast a Gaussian probability distribution for each latent DOF. To ensure the efficiency and accuracy of GP’s predictions, we provide additional accuracy measures. First, we define prediction of a particular DOF as healthy, if the corresponding true solution lies within two standard deviations of the mean value.
@@ 3.5 GP performance in the latent space
Figure 9: Latent distribution prediction for two randomly chosen test examples from the liver test dataset. All the 16 true latent values are transformed to an uniform scale, along with the displacement random variables. True latents lying within the two standard deviations (±$\pm 2\sigma$ region represented by shaded gray region) are indicated with green lines, while the latents lying outside are indicated with red colors. (a) All true latent displacements lie within ±$\pm 2\sigma$, which we classify as a correct prediction. (b) Three true latent values don’t lie within ±$\pm 2\sigma$, which we classify as an incorrect prediction.
@@ 3.6 Uncertainty for missing data region
In the second case, we use a new test dataset in which two dimensional input forces are randomly generated as denoted by red dots in Figure 10b. In Figure 11b we plot the uncertainty for the first latent DOF. It is evident that the uncertainty shows an increasing trend in the region not supported by training data. All the 20 latent DOFs show a very similar trend. In the presented examples, these uncertainties correlate with the GP prediction errors, which can provide useful information. However, it’s crucial to emphasize that the predictive capabilities in the data-sparse and extrapolated regions can never be assured, with the tendency to deteriorate with the distance from the training data.
@@
Similarly, higher uncertainty values in regions not supported by data are also observed in the full field space. For instance, one of the test points located in the white region of Figure 10b has its prediction projected into the full space. As shown in Figure 11b, the latent uncertainty for such an input is high. Similarly, Figure 12f shows that the predicted full field uncertainties are also high throughout the 2D domain. As mentioned earlier, uncertainties are only responsible for tracking the GP component error and are not directly correlated with the autoencoder errors. However, higher levels of uncertainty values, which exceed acceptable engineering thresholds, serve as a reliable indicator of whether the full field solution should be trusted or not.
(a) Framework solution
(b) FEM solution
(c) Absolute error, $e_{\text{f}}$, see Eq. (20)
(d) Reconstruction error, $e_{r}$, see Eq. (21)
(e) GP component error, $e_{\text{GP}}$, see Eq. (22)
(f) Uncertainty prediction, $(u_{\sigma}^{*})^{2}$, see Eq. (14)
Figure 12: Deformation of the 2D beam subjected to a point load (-0.14, 0.15) N on the corner node, which is present in the missing training data region as described by white circle in Figure 10b. Nodal displacements obtained using the (a) proposed framework (b) FEM respectively. (c) Nodal error and (d) reconstruction error of the framework. (e) GP component error. (f) Uncertainty predicted by the framework for the full field solution.
@@ 4 Conclusions
This work presents a novel approach to probabilistic surrogate modeling of high-fidelity simulations. It combines Gaussian processes with reduced-order modeling to efficiently simulate the mechanics of solids while accounting for uncertainties. The reduced representations of high-dimensional displacement data are achieved using autoencoder neural networks, allowing for drastic non-linear compression. For instance, this method reduced approximately 9100 degrees of freedom (DOFs) to 16 latent-space components, as demonstrated with the liver example. This strategy significantly reduces the computational cost of probabilistic predictions using Gaussian processes by lowering the output dimensionality, which is crucial because this cost can scale linearly to cubically with the output dimension size. We demonstrate that the proposed framework can accurately predict non-linear hyper-elastic deformations of solid bodies, along with the associated model uncertainties.
@@
This study opens up several other opportunities for future research. One exciting direction is applying this approach to probabilistic simulations of time and path dependent problems. One could use latent state to track how solutions evolve in such cases, which would make the whole process much faster and more efficient [67]. Another direction would be to exploit state of the art graph autoencoder networks to find compressed representations for arbitrary high-dimensional meshes [68]. This advancement would overcome the limitations of fully connected networks, which are recognized for their challenges in handling problems with a high number of dimensions. Finally, we believe that the proposed approach can find direct application in research and development, which is not restricted to problems in mechanics but covers a broader spectrum of engineering and scientific domains. To facilitate this, we have made all the codes and datasets used in this work available as open-access in the GitHub repository at https://github.com/saurabhdeshpande93/gp-auto-regression.
@@
Acknowledgements:
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No. 764644. Jakub Lengiewicz would like to acknowledge the support from EU Horizon 2020 Marie Sklodowska Curie Individual Fellowship MOrPhEM under Grant 800150. Jakub Lengiewicz would like to acknowledge the support of the Luxembourg National Research Fund (FNR) through the project SUMO, Grant INTER/GACR/21/16555380. Stéphane Bordas and Jakub Lengiewicz are grateful for the support of the Fonds National de la Recherche Luxembourg FNR grant QuaC C20/MS/14782078. Stéphane Bordas received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 811099 TWINNING Project DRIVEN for the University of Luxembourg. This paper only contains the author’s views and the Research Executive Agency and the Commission are not responsible for any use that may be made of the information it contains.